Two identical uniform spheres A and B, with mass M = 1. 00 [kg], are

placed 0. 500 [m] away from another spherical mass C (m =

0. 300 [kg]) as shown. If mass C is released from rest, what is its net

acceleration? Assume that only the gravitational forces from spheres A

and B act on C.


1
Expert's answer
2021-06-02T09:36:40-0400


Mass of identical sphere(M)=1 kg

Mass of another sphere (m)=0.3 kg

Distance between sphere(d)=0.50 m

From the above figure,

Let us calculate the angle C by the law of cosines:

"cosC=\frac{AC^2+BC^2-AB^2}{2 \times AC \times BC} \\

= \frac{0.5^2+0.5^2-0.8^2}{2 \times 0.5\times 0.5} = -0.28"

Force due to A on C is given as:

"F_{AC}=\frac{Gm_Am_C}{AC^2}=\frac{GMm}{0.5^2}=F"

Force due to B on C is given as:

"F_{BC}=\frac{Gm_Am_C}{BC^2}= \frac{GMm}{0.5^2}=F"

The resultant or net force on C is given as:

"F_{net}= \sqrt{F_{AC}^2+F_{BC}^2+2F_{AC}F_{BC}cosC} \\

F_{AC}=F_{BC}=F \\

F_{net}=F\sqrt{2+2cosC} \\

=\frac{GMm}{d^2}\sqrt{2+2cosC} \\

= \frac{GMm}{0.5^2}\sqrt{2+2(-0.28)} = 4.8 \;Gmm"

Acceleration

"a= \frac{F_{net}}{m} \\

= \frac{4.8 \;Gmm}{m}= 4.8\;GM"

G = gravitational constant

"G= 6.67 \times 10^{-11} \;N m^2/kg^2 \\

a = 4.8 \times 6.67 \times 10^{-11} \times 1 = 32.03 \times 10^{-11} \;m/s^2"

Answer: "32.03 \times 10^{-11} \;m/s^2"